OpenStudy (anonymous):

Calculate the derivative of e^(-t) * cos(t) at the point t=(pie/2)

OpenStudy (anonymous):

f*g=f'*g+f*g' therefore, f = e^(-t), g=cos(t). (f will use chain rule)

OpenStudy (anonymous):

then substitute pi/2 into the resulting expression to find the instantaneous speed.